OpenStudy (anonymous):

How do you solve system of equations 6x+4y=32 -6x+4y=8

OpenStudy (anonymous):

hey you need some help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

LOL BTW..waht grade r u in..

OpenStudy (anonymous):

10th lol i feel stupid bc i dnt kno it, its 4 apex

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

so wnna help me

OpenStudy (anonymous):

anyways srry

OpenStudy (anonymous):

im so srry but i GTG

OpenStudy (anonymous):

use sum to product formula , watever its called 6x+4y=32 -6x+4y=8 ----------- 0x +8y=40 y=5 6x +4(5)=32 6x+20=32 6x=32-20 6x=12 x=2

OpenStudy (anonymous):

shanxs

OpenStudy (anonymous):

np